# SALSA: Swift Adaptive Lightweight Self-Attention for Enhanced LiDAR Place Recognition

### v1

@@ Abstract
Large-scale LiDAR mappings and localization leverage place recognition techniques to mitigate odometry drifts, ensuring accurate mapping. These techniques utilize scene representations from LiDAR point clouds to identify previously visited sites within a database. Local descriptors, assigned to each point within a point cloud, are aggregated to form a scene representation for the point cloud. These descriptors are also used to re-rank the retrieved point clouds based on geometric fitness scores. We propose SALSA, a novel, lightweight, and efficient framework for LiDAR place recognition. It consists of a Sphereformer backbone that uses radial window attention to enable information aggregation for sparse distant points, an adaptive self-attention layer to pool local descriptors into tokens, and a multi-layer-perceptron Mixer layer for aggregating the tokens to generate a scene descriptor. The proposed framework outperforms existing methods on various LiDAR place recognition datasets in terms of both retrieval and metric localization while operating in real-time.
@@ I INTRODUCTION
Traditional LiDAR place recognition (LPR) systems use hand-crafted statistics and histogram-based approaches [5, 6, 7], which make them unsuitable for describing large and complex scenes. With the introduction of large-scale datasets [8, 9, 10, 11], several deep learning based solutions have been proposed [1, 2, 3, 4]. These methods typically consist of a deep learning network that extracts descriptors for each point in a point cloud and an aggregator module to combine these features into a unified embedding. Despite significant advancements, existing state-of-the-art methods struggle to achieve accurate localization while being computationally efficient and lightweight. Thus, we propose SALSA which improves localization and retrieval performance on six large-scale datasets while being memory and compute efficient.
@@ IV-A Datasets and Evaluation Methodology
IV-B Results
IV-B1 Retrieval Results
The retrieval performance on ‘easy’ and ‘hard’ sets without re-ranking is detailed in Tab. I and II, respectively. We used the same metrics as in [14] to ensure a consistent evaluation and compared our results with those of Locus [20], DISCO [35], DH3D [35], MinkLoc3D [29], MinkLoc3Dv2 [30], PPT-Net [28], EgoNN [2], LCDNet [3], and LoGG3D-Net [4]. These results demonstrate the efficacy and superior performance of SALSA for place recognition. Although LoGG3D-Net achieves a marginally higher Recall@5 on the ‘easy’ set, SALSA outperforms it in all other metrics on average.
@@ IV-B2 Re-ranking Results
The re-ranking performance evaluation of SALSA with other recent methods is shown in Tab. III. We apply the re-ranking proposed in SpectralGV [14] on the local descriptors of EgoNN [2], LCDNet [3], LoGG3D-Net [4], and SALSA (ours) to rerank the top 20 candidates for each query similar to evaluation proposed in SpectralGV [14].
@@ IV-B4 Discussion of Results
Performance in ‘hard’ set: This set features shorter, unseen sequences with larger gaps between LiDAR scans, leading to a general performance decline across all methods. Tab. I and II: While LoGG3D-Net and EgoNN slightly outperform SALSA on KITTI and Mulran, respectively, they underperform significantly on datasets like ALITA, where the data collection design markedly differs from the ones in the training data. Conversely, SALSA generalizes well to all datasets and obtains the best results on average, demonstrating its superior adaptiveness and resistance to overfitting. Post-re-ranking performance: The performance of LCDNet after re-ranking increases by a higher percentage as compared to the other methods due to its large 640-dimension local descriptor, sacrificing inference speed (Fig. 7). This necessitates the downsampling of points before aggregation, leading to mediocre outcomes in Tab.  I and II. In contrast, SALSA maintains high performance and inference speed with a smaller 16-dimension descriptor, both before and after re-ranking, also illustrating its local descriptors’ robust descriptiveness.
@@ IV-F4 Random Rotation/Occlusion
Evaluating SALSA with dynamic rotation of up to 30 degrees around the z-axis and removal of points in random 30-degree sectors yielded average $F1_{max}$ scores of 0.83 in both cases, demonstrating the consistent performance of SALSA and its robustness to rotation and occlusion.

### v2

@@ Abstract
Large-scale LiDAR mappings and localization leverage place recognition techniques to mitigate odometry drifts, ensuring accurate mapping. These techniques utilize scene representations from LiDAR point clouds to identify previously visited sites within a database. Local descriptors, assigned to each point within a point cloud, are aggregated to form a scene representation for the point cloud. These descriptors are also used to re-rank the retrieved point clouds based on geometric fitness scores. We propose SALSA, a novel, lightweight, and efficient framework for LiDAR place recognition. It consists of a Sphereformer backbone that uses radial window attention to enable information aggregation for sparse distant points, an adaptive self-attention layer to pool local descriptors into tokens, and a multi-layer-perceptron Mixer layer for aggregating the tokens to generate a scene descriptor. The proposed framework outperforms existing methods on various LiDAR place recognition datasets in terms of both retrieval and metric localization while operating in real-time. The code is available at https://github.com/raktimgg/SALSA.
@@ I INTRODUCTION
Traditional LiDAR place recognition (LPR) systems use hand-crafted statistics and histogram-based approaches [5, 6, 7, 8], which make them unsuitable for describing large and complex scenes. With the introduction of large-scale datasets [9, 10, 11, 12], several deep learning based solutions have been proposed [2, 3, 1, 4]. These methods typically consist of a deep learning network that extracts descriptors for each point in a point cloud and an aggregator module to combine these features into a unified embedding. Despite significant advancements, existing state-of-the-art methods struggle to achieve accurate localization while being computationally efficient and lightweight. Thus, we propose SALSA which improves localization and retrieval performance on six large-scale datasets while being memory and compute efficient.
@@ IV-A Datasets and Evaluation Methodology
IV-B Results
IV-B1 Retrieval Results
The retrieval performance on ‘easy’ and ‘hard’ sets without re-ranking is detailed in Tab. I and II, respectively, using the same evaluation metrics as in [15] for consistency. We compare with Locus [21], DISCO [36], DH3D [36], MinkLoc3D [30], MinkLoc3Dv2 [31], PPT-Net [29], EgoNN [3], LCDNet [1], and LoGG3D-Net [4]. These results demonstrate the efficacy and superior performance of SALSA for place recognition. Although LoGG3D-Net achieves a marginally higher Recall@5 on the ‘easy’ set, SALSA outperforms it in all other metrics on average.
@@ IV-B2 Re-ranking Results
The re-ranking performance evaluation of SALSA and other recent methods is shown in Tab. III. We apply the re-ranking method from SpectralGV [15] to the local descriptors of EgoNN [3], LCDNet [1], LoGG3D-Net [4], and SALSA (ours), re-ranking the top 20 candidates for each query as in SpectralGV [15].
@@ IV-B4 Discussion of Results
Performance in ‘hard’ set: This set features shorter, unseen sequences with larger gaps between LiDAR scans, leading to a general performance decline across all methods. Tab. I and II: While LoGG3D-Net and EgoNN slightly outperform SALSA on KITTI and Mulran, respectively, they underperform significantly on datasets like ALITA, where the data collection design markedly differs from the ones in the training data. Conversely, SALSA generalizes well to all datasets and obtains the best results on average, demonstrating its superior adaptiveness and resistance to overfitting. Post-re-ranking performance: The performance of LCDNet after re-ranking increases by a higher percentage as compared to the other methods due to its large 640-dimension local descriptor, sacrificing inference speed (Fig. 7). This necessitates downsampling points before aggregation, leading to mediocre outcomes in Tab.  I and II. In contrast, SALSA maintains high performance and speed with a smaller 16-dimensional descriptor, demonstrating robust local descriptors both before and after re-ranking.
@@ IV-F4 Random Rotation/Occlusion
Evaluating SALSA with dynamic rotation of up to 30 degrees around the z-axis and removal of points in random 30-degree sectors yielded average $F1_{max}$ scores of 0.83 in both cases, demonstrating SALSA’s robustness to rotation and occlusion.